Two long, parallel wires are separated by 7.89 cm and carry currents of 2.23 A and 3.93 A, respectively. Find the magnitude of the magnetic force ? that acts on a 2.23 m length of either wire.

F = ?N
Answer #1

here,

the seperation between the wires , d = 7.89 cm = 0.0789 m

current in wire 1 , I1 = 2.23 A

current in wire 2 , I2 = 3.93 A

length of wire , L = 2.23 m

the magnitude of the magnetic force that acts on 2.23 m length of either wire , F = u0 * I1 * I2 * L /( 2*pi*d)

F = 1.26 * 10^-6 * 2.23 * 3.93 * 2.23 /(2*pi*0.0789) N

F = 4.97 * 10^-5 N

the magnitude of the magnetic force that acts on 2.23 m length of either wire is 4.97 * 10^-5 N